OpenStudy (anonymous):

Find the value(s) of x that make the following equation true: √-4x+5 =x

OpenStudy (anonymous):

square both sides

OpenStudy (anonymous):

and then you solve the quadratic

OpenStudy (anonymous):

-4x+5 = x^2

OpenStudy (anonymous):

0 = x^2+4x-5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then.....

OpenStudy (anonymous):

then factorize the quadratic.

OpenStudy (anonymous):

Ya, I'm having trouble with that part.

OpenStudy (anonymous):

we need two numbers such that (a) their product = -5 (b) their sum = +4

OpenStudy (anonymous):

can you think of any?

OpenStudy (anonymous):

-1 and 5

OpenStudy (anonymous):

5 and -1

OpenStudy (anonymous):

good no, \[0=(x-1)(x+5)\]

OpenStudy (anonymous):

so, \[x=1\qquad{\rm AND}\qquad x=-5 \]

OpenStudy (anonymous):

Ya, i was getting to that. Thanks

OpenStudy (anonymous):

lets see if they both satisfy the given problem or not

OpenStudy (anonymous):

plug in x=1 and see if it satisfies the given equation before you squared it!

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now, what about x= -5 ??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

good coz, \[\sqrt{-4(-5)+5}=\sqrt{25}=5\ne-5\]

OpenStudy (anonymous):

so, there is only one value for "x" that can satisfy the given equation and that is, x=1

OpenStudy (anonymous):

yes. Thanks

OpenStudy (anonymous):

good job people.